03 January 2008

RHIC Startup:  Getting going again after the maintenance day was not smooth.  Retuning of both beams was necessitated by the polarity flip of the PHENIX magnets.  Pressure in the PHENIX straight section bi8-cc-pw3.1 increases by nearly 3 orders of magnitude.  The PHENIX IR vacuum shows no pressure bump.  After getting a 6x6 bunch store in around 10:00PM last night they attempted an 87x87 bunch store.  Yellow efficiency is down, and Blue exhibits a drop when Yellow is injected. 

STAR was unable to work on their water leak because of hydraulic problems in the pole tip extraction. 

PHENIX did the magnetic field flip, but were unable to fix the Pad Chamber DAQ problem. 

Baby bunches in AGS.  When Peggy is able to resolve this problem, the experiments request that the beam be dumped and refilled. 

Booster Main Magnet Power Supply (MMPS) problem is keeping us from refilling in the morning. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020